OpenStudy (anonymous):

Solve the equation: \[2x ^{1/2}=18\]

OpenStudy (apoorvk):

square both sides. you get. \[4x = 18^2\]\[or, x = \frac{18\times 18}{4}\]

OpenStudy (apoorvk):

@Eyad - I think you made a small wee bit of mistake. :P

OpenStudy (anonymous):

LOL @apoorvk :I think u didn't read the question clearly :P

OpenStudy (anonymous):

Depend if the power on x or (2x) ..

OpenStudy (apoorvk):

@Eyad , it's: \[2 \sqrt x = 18\]

OpenStudy (apoorvk):

well, there are no brackets in the question posted above, so mine is definitely (or accidentally :P) right in here it seems.

OpenStudy (apoorvk):

@mikebmxer - do read this up again.

OpenStudy (anonymous):

Well ,angry birds should win That Boy <<<< Useless i thought he gonna bring Luck !

OpenStudy (apoorvk):

Looooollll!!! :D

OpenStudy (anonymous):

xD